OpenStudy (anonymous):

HELP!!!!! 3^2x+1=1/81

OpenStudy (jtvatsim):

this one is actually pretty easy with a few tricks.

OpenStudy (jtvatsim):

\[3^{2x+1} = \frac{1}{81}\] \[3^{2x+1} = \frac{1}{3^4}\] \[3^{2x+1} = 3^{-4}\]

OpenStudy (anonymous):

then I can cancel the 3s

OpenStudy (jtvatsim):

yes, technically you are taking log(base 3) both sides :)

OpenStudy (anonymous):

then its 2x+1=-4

OpenStudy (jtvatsim):

exactly!

OpenStudy (anonymous):

sweet then solve from there!!

OpenStudy (jtvatsim):

perfect, you've got it!

OpenStudy (anonymous):

answer is

OpenStudy (anonymous):

x=-5/2?

OpenStudy (jtvatsim):

let's see....... yes!

OpenStudy (anonymous):

perfect ty

OpenStudy (jtvatsim):

yw